# Information-Theoretic Bayesian Optimization for Bilevel Optimization Problems

### v1

@@ 3.1 Lower Bound of Mutual Information
$p(y^{f}_{(\boldsymbol{x},\boldsymbol{\theta})}\mid f(\boldsymbol{x},\boldsymbol{\theta}^{*}(\boldsymbol{x}))\leq f^{*},{\cal D}_{t}^{+})$ $=\begin{cases}\Phi\left(\frac{f^{*}-m_{1}^{f}}{s_{1}^{f}}\right)\phi\left(\frac{y^{f}_{(\boldsymbol{x},\boldsymbol{\theta})}-m_{3}^{f}}{s_{3}^{f}}\right)/\left\{\Phi\left(\frac{f^{*}-m_{2}^{f}}{s_{2}^{f}}\right)s_{3}^{f}\right\}&\text{ if }\boldsymbol{x}\neq\boldsymbol{x}^{*},\\ \phi\left(\frac{y^{f}_{(\boldsymbol{x},\boldsymbol{\theta})}-m_{3}^{f}}{s_{3}^{f}}\right)/s_{3}^{f}&\text{ otherwise, }\end{cases}$ (6)
@@
The proof is in Appendix A.1. Note that all of $\{(m_{i}^{f},s_{i}^{f})\}_{i=1}^{3}$ can be analytically calculated from the GP posterior of $f$ for which details are also show in Appendix A.1. For the numerator in the second term of (5) can be reduced to the similar analytical form, which is shown in Appendix A.2. As a result, we obtain an analytical form of the inside of the expectation (5).
@@ 3.2 Computations
The expectation of (5) is approximated by the Monte-Carlo method by sampling $\Omega\coloneqq\{y^{f}_{(\boldsymbol{x},\boldsymbol{\theta})},y^{g}_{(\boldsymbol{x},\boldsymbol{\theta})},f^{*},g^{*},\boldsymbol{x}^{*},\boldsymbol{\theta}^{*}\}$. The sample of all the elements of $\Omega$ can be obtained through the sample of the objective functions $f$ and $g$. We use random Fourier feature (RFF) (Rahimi and Recht,, 2008), by which the GP posterior can be approximated by the Bayesian linear model. For a $D$-dimensional RFF vector $\boldsymbol{\phi}(\boldsymbol{x},\boldsymbol{\theta})\in\mathbb{R}^{D}$, the linear model $\boldsymbol{w}^{h\top}\boldsymbol{\phi}(\boldsymbol{x},\boldsymbol{\theta})$ can be constructed, where $\boldsymbol{w}^{h}\in\mathbb{R}^{D}$ is a parameter vector and $h\in\{f,g\}$ represents one of objective functions. By sampling $\boldsymbol{w}^{h}$ from the posterior, approximate sample paths of $f$ and $g$ are obtained, which denoted as $\tilde{f}$ and $\tilde{g}$, respectively. Then, the sample of $(f^{*},g^{*},\boldsymbol{x}^{*},\boldsymbol{\theta}^{*})$ is obtained through $\max_{\boldsymbol{x}}\tilde{f}(\boldsymbol{x},\tilde{\boldsymbol{\theta}}^{*}(\boldsymbol{x}))\text{ s.t. }\tilde{\boldsymbol{\theta}}^{*}(\boldsymbol{x})=\arg\max_{\boldsymbol{\theta}}\tilde{g}(\boldsymbol{x},\boldsymbol{\theta})$, which can be seen as a white-box bilevel optimization problem. Since both $\tilde{f}$ and $\tilde{g}$ are represented by the Bayesian liner model, they are differentiable. Then, the gradient $\partial\tilde{f}(\boldsymbol{x},\boldsymbol{\theta}^{*}(\boldsymbol{x}))/\partial\boldsymbol{x}$ can be obtained through the implicit function theorem (see Appendix B for detail), by which standard gradient based optimization methods can be applied. The samples of $y^{f}_{(\boldsymbol{x},\boldsymbol{\theta})}$ and $y^{g}_{(\boldsymbol{x},\boldsymbol{\theta})}$ can be obtained by adding the random noise from ${\cal N}(0,\{\sigma^{f}_{\rm noise}\}^{2})$ and ${\cal N}(0,\{\sigma^{g}_{\rm noise}\}^{2})$ to the sampled $f(\boldsymbol{x},\boldsymbol{\theta})$ and $g(\boldsymbol{x},\boldsymbol{\theta})$, respectively.
@@ 4.1 Decoupled Setting
$\mathrm{MI}(y^{f}_{(\boldsymbol{x},\boldsymbol{\theta})}\ ;\ f^{*},g^{*},\boldsymbol{x}^{*},\boldsymbol{\theta}^{*}\mid{\cal D}_{t})$ $\geq\mathbb{E}_{\Omega}\left[\log\frac{p(y^{f}_{(\boldsymbol{x},\boldsymbol{\theta})}\mid f(\boldsymbol{x},\boldsymbol{\theta}^{*}(\boldsymbol{x}))\leq f^{*},{\cal D}_{t}^{+})}{p(y^{f}_{(\boldsymbol{x},\boldsymbol{\theta})}\mid{\cal D}_{t})}\right],$ (7)
$\mathrm{MI}(y^{g}_{(\boldsymbol{x},\boldsymbol{\theta})}\ ;\ f^{*},g^{*},\boldsymbol{x}^{*},\boldsymbol{\theta}^{*}\mid{\cal D}_{t})$ $\geq\mathbb{E}_{\Omega}\left[\log\frac{p(y^{g}_{(\boldsymbol{x},\boldsymbol{\theta})}\mid g(\boldsymbol{x}^{*},\boldsymbol{\theta})\leq g^{*},{\cal D}_{t}^{+})}{p(y^{g}_{(\boldsymbol{x},\boldsymbol{\theta})}\mid{\cal D}_{t})}\right].$ (8)
@@ 4.2 Incorporating Constraint Problems
(c) Energy
(d) SMD01
(e) SMD02
(f) SMD03
Figure 3: Regret comparison on benchmark problems.
@@ 5 Experiments
$r_{f}(\boldsymbol{x}_{i},\boldsymbol{\theta}_{i})$ $=\max(0,f^{*}-f(\boldsymbol{x}_{i},\boldsymbol{\theta}_{i}))/(f^{*}-\min_{\boldsymbol{x},\boldsymbol{\theta}}f(\boldsymbol{x},\boldsymbol{\theta})),$
$r_{g}(\boldsymbol{x}_{i},\boldsymbol{\theta}_{i})$ $=(g(\boldsymbol{x}_{i},\boldsymbol{\theta}^{*}(\boldsymbol{x}_{i}))-g(\boldsymbol{x}_{i},\boldsymbol{\theta}_{i}))/(g(\boldsymbol{x}_{i},\boldsymbol{\theta}^{*}(\boldsymbol{x}_{i}))-\min_{\boldsymbol{\theta}}g(\boldsymbol{x}_{i},\boldsymbol{\theta})).$
@@
We first provide the results on coupled setting for GP sample path functions (section 5.1) and benchmark functions (section 5.2). Further, the results on decoupled setting (section 5.3) and different $K$ settings (section 5.4) are also reported. Appendix presents other details of the settings (Appendix E.1) and results such as a larger noise setting (Appendix E.2), the continuous domain (Appendix E.5), and constraint problems (Appendix E.6).
@@ 5.1 Sample Path from GP Prior
5.2 Benchmark Functions
We here used six benchmark problems. Two functions are created by combining benchmark functions of single-level optimization. In the first problem, denoted as BG, the upper objective is BraninHoo ($d_{{\cal X}}=1$) and the lower objective Goldstein-price ($d_{\Theta}=1$), which was used in (Chew et al.,, 2025). In the second problem, denoted as SB, the upper objective is SixHumpCamel ($d_{{\cal X}}=1$) and the lower objective BraninHoo ($d_{\Theta}=1$), which was used in (Ekmekcioglu et al.,, 2024). The third problem, denoted as Energy, is a simulator based energy market problem ($d_{{\cal X}}=2$ and $d_{\Theta}=2$) introduced by (Chew et al.,, 2025), in which this data is regarded as a real-world dataset. From the fourth to the sixth problems, denoted as SMD01, 02, and 03 ($d_{{\cal X}}=2$ and $d_{\Theta}=2$), are test problems specifically designed for bilevel optimization benchmark (Sinha et al.,, 2014). The number of grid points in each dimension is $100$ for GB and SB ($100^{2}$ points), and $10$ for Energy and SMD ($10^{4}$ points).
The results are shown in Fig. 3. BLJES has obviously superior performance in BG, SB, Energy, and SMD02. For SMD01 and SMD03, similar performance is shown in BLJES and BILBO, both of which rapidly decrease the regret compared with Random.
@@ 5.3 Decoupled Setting
(e) SMD01
(f) SMD02
(g) SMD03
Figure 4: Regret comparison on benchmark problems under decoupled setting.
@@ 6 Conclusion
We propose an information-theoretic approach to bilevel Bayesian optimization, called Bilevel optimization via Lower-bound based Joint Entropy Search (BLJES). BLJES considers information gain of optimal points and values of both the upper- and lower- level problems simultaneously, by which we can define a unified criterion that measures the benefit for both the problems. We derive a lower bound based approximation of bilevel information gain, which can be seen as a natural extension of the single level information-theoretic Bayesian optimization. Further, we also propose extensions for decoupled setting and constraint problems. The effectiveness of BLJES is demonstrated through sample path functions from Gaussian processes and benchmark functions.
Acknowledgement
This work is partially supported by JSPS KAKENHI (Grant Number 23K21696, 25K03182, 22H00300, and 23K17817), and Data Creation and Utilization Type Material Research and Development Project (Grant No. JPMXP1122712807) of MEXT.
@@ E.1 Other Details of Experimental Settings
We used the SingleTaskGP model of BoTorch (Balandat et al.,, 2020) to define the GPs. The output of each benchmark function are transformed by signed log1p function ${\rm sign}(y)\log(1+|y|)$ except for BraninHoo and Goldstein-price for which the transformation shown by (Picheny et al.,, 2013) was used. For benchmark functions, the input space is scaled to $0,1]^{d_{{\cal X}}+d_{\Theta}}$ from the original input domain.
@@ E.3 Additional Results on Decoupled Setting
Figure 7 shows regret in decoupled setting for all different length scale settings of the GP prior, which generates true objectives. We obviously see that BLJES was superior or comparable to BILBO.

### v2

@@ 3.1 Lower Bound of Mutual Information
$p(y^{f}_{(\boldsymbol{x},\boldsymbol{\theta})}\mid f(\boldsymbol{x},\boldsymbol{\theta}^{*}(\boldsymbol{x}))\leq f^{*},{\cal D}_{t}^{+})$
$=\begin{cases}\frac{\Phi\left(\frac{f^{*}-m_{1}^{f}}{s_{1}^{f}}\right)\phi\left(\frac{y^{f}_{(\boldsymbol{x},\boldsymbol{\theta})}-m_{3}^{f}}{s_{3}^{f}}\right)}{\Phi\left(\frac{f^{*}-m_{2}^{f}}{s_{2}^{f}}\right)s_{3}^{f}}&\text{ if }\boldsymbol{x}\neq\boldsymbol{x}^{*},\\ \phi\left(\frac{y^{f}_{(\boldsymbol{x},\boldsymbol{\theta})}-m_{3}^{f}}{s_{3}^{f}}\right)/s_{3}^{f}&\text{ otherwise, }\end{cases}$ (6)
@@
The proof is in Appendix A.1. Note that all of $\{(m_{i}^{f},s_{i}^{f})\}_{i=1}^{3}$ can be analytically calculated from the GP posterior of $f$ easily, for which details are also show in Appendix A.1. For the numerator in the second term of (5) can be reduced to the similar analytical form, which is shown in Appendix A.2. As a result, we obtain an analytical form of the inside of the expectation (5).
@@ 3.2 Computations
The expectation of (5) is approximated by the Monte-Carlo method by sampling $\Omega\coloneqq\{y^{f}_{(\boldsymbol{x},\boldsymbol{\theta})},y^{g}_{(\boldsymbol{x},\boldsymbol{\theta})},f^{*},g^{*},\boldsymbol{x}^{*},\boldsymbol{\theta}^{*}\}$. The sample of all the elements of $\Omega$ can be obtained through the sample of the objective functions $f$ and $g$. We use random Fourier feature (RFF) (Rahimi and Recht,, 2008), by which the GP posterior can be approximated by the Bayesian linear model. For a $D$-dimensional RFF vector $\boldsymbol{\phi}(\boldsymbol{x},\boldsymbol{\theta})\in\mathbb{R}^{D}$, the linear model $\boldsymbol{w}^{h\top}\boldsymbol{\phi}(\boldsymbol{x},\boldsymbol{\theta})$ can be constructed, where $\boldsymbol{w}^{h}\in\mathbb{R}^{D}$ is a parameter vector and $h\in\{f,g\}$ represents one of objective functions. By sampling $\boldsymbol{w}^{h}$ from the posterior, approximate sample paths of $f$ and $g$ are obtained, which are denoted as $\tilde{f}$ and $\tilde{g}$, respectively. Then, the sample of $(f^{*},g^{*},\boldsymbol{x}^{*},\boldsymbol{\theta}^{*})$ is obtained through $\max_{\boldsymbol{x}}\tilde{f}(\boldsymbol{x},\tilde{\boldsymbol{\theta}}^{*}(\boldsymbol{x}))\text{ s.t. }\tilde{\boldsymbol{\theta}}^{*}(\boldsymbol{x})=\arg\max_{\boldsymbol{\theta}}\tilde{g}(\boldsymbol{x},\boldsymbol{\theta})$, which can be seen as a white-box bilevel optimization problem. Since both $\tilde{f}$ and $\tilde{g}$ are represented by the Bayesian liner model, they are differentiable. Then, the gradient $\partial\tilde{f}(\boldsymbol{x},\boldsymbol{\theta}^{*}(\boldsymbol{x}))/\partial\boldsymbol{x}$ can be obtained through the implicit function theorem (see Appendix B for detail), by which standard gradient based optimization methods can be applied. The samples of $y^{f}_{(\boldsymbol{x},\boldsymbol{\theta})}$ and $y^{g}_{(\boldsymbol{x},\boldsymbol{\theta})}$ can be obtained by adding the random noise from ${\cal N}(0,\{\sigma^{f}_{\rm noise}\}^{2})$ and ${\cal N}(0,\{\sigma^{g}_{\rm noise}\}^{2})$ to the sampled $f(\boldsymbol{x},\boldsymbol{\theta})$ and $g(\boldsymbol{x},\boldsymbol{\theta})$, respectively.
@@ 4.1 Decoupled Setting
$\mathrm{MI}(y^{f}_{(\boldsymbol{x},\boldsymbol{\theta})}\ ;\ f^{*},g^{*},\boldsymbol{x}^{*},\boldsymbol{\theta}^{*}\mid{\cal D}_{t})$
$\ \ \geq\mathbb{E}_{\Omega}\left[\log\frac{p(y^{f}_{(\boldsymbol{x},\boldsymbol{\theta})}\mid f(\boldsymbol{x},\boldsymbol{\theta}^{*}(\boldsymbol{x}))\leq f^{*},{\cal D}_{t}^{+})}{p(y^{f}_{(\boldsymbol{x},\boldsymbol{\theta})}\mid{\cal D}_{t})}\right],$ (7)
$\mathrm{MI}(y^{g}_{(\boldsymbol{x},\boldsymbol{\theta})}\ ;\ f^{*},g^{*},\boldsymbol{x}^{*},\boldsymbol{\theta}^{*}\mid{\cal D}_{t})$
$\ \ \geq\mathbb{E}_{\Omega}\left[\log\frac{p(y^{g}_{(\boldsymbol{x},\boldsymbol{\theta})}\mid g(\boldsymbol{x}^{*},\boldsymbol{\theta})\leq g^{*},{\cal D}_{t}^{+})}{p(y^{g}_{(\boldsymbol{x},\boldsymbol{\theta})}\mid{\cal D}_{t})}\right].$ (8)
@@ 4.2 Incorporating Constraint Problems
(c) SMD01
(d) SMD02
(e) SMD03
(f) Energy
(g) Chemical
(h) Material
Figure 4: Regret comparison on benchmark and real-world problems.
@@ 5 Experiments
$r_{f}(\boldsymbol{x}_{i},\boldsymbol{\theta}_{i})$ $=\frac{\max(0,f^{*}-f(\boldsymbol{x}_{i},\boldsymbol{\theta}_{i}))}{(f^{*}-\min_{\boldsymbol{x},\boldsymbol{\theta}}f(\boldsymbol{x},\boldsymbol{\theta}))},$
$r_{g}(\boldsymbol{x}_{i},\boldsymbol{\theta}_{i})$ $=\frac{(g(\boldsymbol{x}_{i},\boldsymbol{\theta}^{*}(\boldsymbol{x}_{i}))-g(\boldsymbol{x}_{i},\boldsymbol{\theta}_{i}))}{(g(\boldsymbol{x}_{i},\boldsymbol{\theta}^{*}(\boldsymbol{x}_{i}))-\min_{\boldsymbol{\theta}}g(\boldsymbol{x}_{i},\boldsymbol{\theta}))}.$
@@
We first provide the results on coupled setting for GP sample path functions (section 5.1) and benchmark and real-world functions (section 5.2). Further, the results on decoupled setting (section 5.3) and different $K$ settings (section 5.4) are also reported. Appendix presents other details of the settings (Appendix E.1) and other empirical evaluations such as a larger noise setting (Appendix E.2), the continuous domain (Appendix E.5), constraint problems (Appendix E.6), higher dimensional settings (Appendix E.7), comparison with a simplified variant of MLJES (Appendix E.8), and the effect of approximations (Appendix E.9).
@@ 5.1 Sample Path from GP Prior
5.2 Benchmark and Real-World Data
We here used the five benchmark and three real-world problems. Two functions are created by combining benchmark functions of single-level optimization. In the first problem, denoted as BG, the upper objective is BraninHoo ($d_{{\cal X}}=1$) and the lower objective Goldstein-price ($d_{\Theta}=1$), which was used in (Chew et al.,, 2025). In the second problem, denoted as SB, the upper objective is SixHumpCamel ($d_{{\cal X}}=1$) and the lower objective BraninHoo ($d_{\Theta}=1$), which was used in (Ekmekcioglu et al.,, 2024). From the third to the fifth problems, denoted as SMD01, 02, and 03 ($d_{{\cal X}}=2$ and $d_{\Theta}=2$), are test problems specifically designed for bilevel optimization benchmark (Sinha et al.,, 2014). The sixth problem, denoted as Energy, is a simulator based energy market problem ($d_{{\cal X}}=2$ and $d_{\Theta}=2$), and the seventh problem, denoted as Chemical, is about an optimization of simulated mass flow of Methyl Acetate ($d_{{\cal X}}=1$ and $d_{\Theta}=3$). Chemical has one constraint function for the upper-level problem. Energy and Chemical data are introduced by (Chew et al.,, 2025), in which these are regarded as a real-world dataset (see Chew et al., (2025) for the detailed definitions). The number of grid points in each dimension is $100$ for GB and SB ($100^{2}$ points), and $10$ for Energy and SMD ($10^{4}$ points). The last problem is Material data (which is newly created by ourselves), consisting of simulation based evaluations for the crystal structures $\text{Fe}_{x}\text{Ni}_{y}\text{Cr}_{z}$. The problem is to maximize the bulk-modulus $f(\boldsymbol{x},\boldsymbol{\theta})$ with respect to $\boldsymbol{x}=(x,y,z)^{\top}$ under the minimization constraint of the energy $g(\boldsymbol{x},\boldsymbol{\theta})$, where $\boldsymbol{\theta}$ corresponds to the three dimensional position of the atoms ($\boldsymbol{\theta}$ itself is so-called structure descriptor whose dimension is $30$). This data have discrete candidates of $54$ upper and $500$ lower variables ($\boldsymbol{x}$ and $\boldsymbol{\theta}$), resulting in 27,000 candidates. See Appendix E.1 for further detail.
The results are shown in Fig. 4. BLJES has obviously superior performance except for SMD01 and SMD03. For SMD01 and SMD03, similar performance is shown in BLJES and BILBO, both of which rapidly decrease the regret compared with Random.
@@ 5.3 Decoupled Setting
Figure 5: Regret comparison on decoupled setting.
@@ 5.4 Effect of the Number of Samplings
6 Limitations
As we described in section 1, the bilevel BO in which both levels are expensive has not been widely studied, and we still have several limitations.
An unresolved theme is the theoretical analysis of the approximation error of MI. The major approximation components of our MI estimator are the lower bound, RFF, and the MC sampling. Providing a theoretical guarantee for the combined approximation error is a challenging issue, though it is actually a common issue for information-theoretic BO methods. For a connection of our lower bound and an independence assumption is discussed in Appendix F, but revealing the approximation error is still an open problem.
Particularly for GP-UCB (Srinivas et al.,, 2010) based approaches including BILBO, convergence of the regret bound has been widely studied. On the other hand, for information-theoretic BO in general, the regret analysis is also still an open problem even for the simplest single-level standard problem setting (see Appendix G). Therefore, the regret analysis for information-theoretic BO including BLJES is still needed to be addressed.
@@ 7 Conclusion
We propose an information-theoretic approach to bilevel Bayesian optimization, called Bilevel optimization via Lower-bound based Joint Entropy Search (BLJES). BLJES considers information gain of optimal points and values of both the upper- and lower-level problems simultaneously, by which we can define a unified criterion that measures the benefit for both the problems. We derive a lower bound based approximation of bilevel information gain, which can be seen as a natural extension of the single level information-theoretic Bayesian optimization. We further propose extensions for decoupled setting and constraint problems. The effectiveness of BLJES is demonstrated through synthetic, benchmark, and real-world functions.
@@ E.1 Other Details of Experimental Settings
We used the SingleTaskGP model of BoTorch (Balandat et al.,, 2020) to define the GPs. The output of each benchmark function are transformed by signed log1p function ${\rm sign}(y)\log(1+|y|)$ except for BraninHoo and Goldstein-price for which the transformation shown by (Picheny et al.,, 2013) was used. For benchmark functions, the input space is scaled to $0,1]^{d_{{\cal X}}+d_{\Theta}}$ from the original input domain. For the chemical dataset, we re-defined the true objective function values as $f(\boldsymbol{x},\boldsymbol{\theta})=-\log((\max_{\boldsymbol{x}^{\prime},\boldsymbol{\theta}^{\prime}}f_{\rm ori}(\boldsymbol{x}^{\prime},\boldsymbol{\theta}^{\prime})+10^{-4})-f_{\rm ori}(\boldsymbol{x},\boldsymbol{\theta}))$ and $g(\boldsymbol{x},\boldsymbol{\theta})=-\log((\max_{\boldsymbol{x}^{\prime},\boldsymbol{\theta}^{\prime}}g_{\rm ori}(\boldsymbol{x}^{\prime},\boldsymbol{\theta}^{\prime})+10^{-4})-g_{\rm ori}(\boldsymbol{x},\boldsymbol{\theta}))$, where $f_{\rm ori}$ and $g_{\rm ori}$ are original functions in the dataset. This is a log transformation applied to the difference from the maximum value (instead of 0). We employed this transformation because, in this dataset, many values are concentrated on the small scale differences around $\max f_{\rm ori}(\boldsymbol{x},\boldsymbol{\theta})$ and $\max g_{\rm ori}(\boldsymbol{x},\boldsymbol{\theta})$, respectively.
In the Material data, the bcc (body-centered cubic) structure $\text{Fe}_{x}\text{Ni}_{y}\text{Cr}_{z}$ with the constraints $x+y+z=54$ and $x,y,z\geq 15$ is considered. For each candidate $(\boldsymbol{x},\boldsymbol{\theta})$, a classical interatomic potential based structure relaxation is performed to obtain the energy and the bulk-modulus. For a given $\boldsymbol{x}$, although the bcc structure specifies the coordinates at which atoms are located, there remains freedom regarding which atom (Fe or Ni or Cr) occupies each site; we randomly generated 500 patterns of atom assignments for each of $\boldsymbol{x}$. For $\boldsymbol{\theta}$, we employed a histogram of atomic distances (we created $20$ bins in $[1,6]$ Å), also called the radial distribution function (RDF), which is a well-known structural descriptor (e.g., Schütt et al.,, 2014). RDF can be defined for each pair of atoms, and after removing redundant dimensions (dimensions with all zero values), we obtain $30$ dimensional $\boldsymbol{\theta}$. Since this data has a high dimensional input, we employ the LogNormal (${\cal L}{\cal N}$) hyper-prior on the Gaussian kernel length scale $\ell$, which follows (Hvarfner et al.,, 2024):
$\ell\sim{\cal L}{\cal N}(\mu_{0}+\log(d_{{\cal X}}+d_{\Theta})/2,\sigma_{0}),$
where $\mu_{0}=\sqrt{2}$ and $\sigma_{0}=\sqrt{3}$ that are suggestions by (Hvarfner et al.,, 2024). This is a common setting for BILBO and BLJES.
@@ E.3 Additional Results on Decoupled Setting
Figure 8 shows regret in decoupled setting for all different length scale settings of the GP prior, and Fig. 9 shows the results on other benchmark functions. We obviously see that BLJES was superior or comparable to BILBO.
@@
(a) SB
(b) SMD01
(c) SMD03
Figure 9: Regret comparison on additional benchmark problems under decoupled setting.